# Concept Unlearning by Modeling Key Steps of Diffusion Process

### v2

@@ Abstract
Text-to-image diffusion models (T2I DMs), represented by Stable Diffusion, which generate highly realistic images based on textual input, have been widely used. However, their misuse poses serious security risks. While existing concept unlearning methods aim to mitigate these risks, they struggle to balance unlearning effectiveness with generative retainability.To overcome this limitation, we innovatively propose the Key Step Concept Unlearning (KSCU) method, which ingeniously capitalizes on the unique stepwise sampling characteristic inherent in diffusion models during the image generation process. Unlike conventional approaches that treat all denoising steps equally, KSCU strategically focuses on pivotal steps with the most influence over the final outcome by dividing key steps for different concept unlearning tasks and fine-tuning the model only at those steps. This targeted approach reduces the number of parameter updates needed for effective unlearning, while maximizing the retention of the model’s generative capabilities.Through extensive benchmark experiments, we demonstrate that KSCU effectively prevents T2I DMs from generating undesirable images while better retaining the model’s generative capabilities. Our code will be released.
Preprint. Under review.
@@
1 Introduction
In recent years, the advancement of text-to-image (T2I) diffusion models [22, 20, 9, 16, 26, 43] has led to their widespread adoption across various domains [38, 31, 42, 1, 27], including short videos, comics, and illustrations. As a result, AI-generated images have become an integral part of daily life for many individuals. However, large-scale T2I models are trained on extensive datasets [45, 4, 37, 38] that often contain sensitive content, such as unauthorized artistic works [46, 47] or Not Safe For Work (NSFW) material. Consequently, these models can inadvertently regenerate infringing content when prompted with relevant text inputs [3]. This raises significant concerns regarding copyright infringement and social security risks, highlighting the urgent need for effective mitigation strategies.
@@ 1 Introduction
The generation of NSFW content poses a major security risk for mainstream text-to-image diffusion models. Researchers have explored various approaches to mitigate this issue, including dataset-based screening, content filtering [39], and fine-tuning techniques [7, 30, 41]. concept unlearning, as a fine-tuning-based approach [23, 11, 56, 13, 14, 53, 18], selectively adjusts partial model parameters to remove the model’s ability to generate specific concepts. Through concept unlearning, T2I diffusion models effectively ‘forget’ designated visual concepts by replacing them with semantically neutral content. This targeted parameter adjustment (referred to as ”concept erase” [12, 21, 14]) provides distinct advantages over conventional methods—offering greater robustness than post-hoc filtering [39] while avoiding the intensive labor of dataset curation [40], and has become one of the most promising approaches to curb the generation of NSFW content in T2I propagation models.
Although significant progress has been made in concept unlearning for T2I diffusion models, several challenges remain unresolved. a. Blind unlearning of the entire diffusion process. Existing concept unlearning approaches for T2I diffusion models fail to notice for the stepwise denoising characteristics of diffusion models. Since diffusion models generate images through a progressive denoising process—where earlier steps primarily capture low-frequency structural information and later steps refine high-frequency details [17, 50, 55, 35, 2]—unlearning does not necessarily require fine-tuning across all denoising steps. b. Unbalanced unlearning effectiveness and generative retainability. Most existing methods require extensive modifications to the original model parameters to ensure unlearning effectiveness. However, excessive parameter updates may lead to over-unlearning, negatively affecting the model’s ability to generate unrelated concepts. Conversely, strictly limiting the scale of parameter updates to preserve generative retainability—i.e., the model’s ability to retain generation quality for unrelated concepts—makes it difficult to fully unlearn the target concept. c. Robustness at high cost. Adversarial attack [51, 6, 58] research on unlearned models has demonstrated that carefully crafted adversarial text prompts can successfully recover unlearned concepts. While adversarial training [21, 59] can enhance robustness against such attacks, it is computationally expensive and brings low generative retainability.
@@
To address existing challenges, we propose Key Step Concept Unlearning (KSCU), a novel framework for concept unlearning in T2I diffusion models. Unlike prior methods that operate across all denoising steps, KSCU selectively fine-tunes a subset of key steps based on a predefined Key Step Table. For different unlearning tasks, multiple step intervals are empirically defined, and a selection algorithm prioritizes steps with the greatest influence on the final output, enabling efficient and targeted unlearning. To improve robustness, KSCU applies Prompt Augmentation in the final training stage, including random shuffling, word removal, character insertion, and embedding noise. These augmented prompts help the model better capture the semantic distribution of the unlearned concept in the embedding space. Additionally, we introduce a Key Step Unlearning Loss tailored to classifier-free guidance DMs. By combining Key Step Table, Prompt Augmentation, and Key Step Unlearning Loss, KSCU achieves more effective concept unlearning while better ensuring generative retainability.
@@
  • We propose KSCU, a novel concept unlearning framework for T2I DMs. Unlike prior work, KSCU only performs fine-tuning for the key step, achieving efficient unlearning by applying smaller adjustments to the model, thereby maintaining the model’s generation ability, taking both unlearning effectiveness and generative retainability into account.
@@
  • Our proposed KSCU integrates three innovations: Key Step Table, a step selection strategy focusing on critical denoising steps to achieve effective concept unlearning with minimal fine-tuning overhead; Prompt Augmentation to enhance robustness by introducing diverse prompt variations; and Key Step Unlearning Loss, a concept unlearning loss tailored for classifier-free guided diffusion models, enabling effective concept removal while preserving generation quality.
  • Extensive experimental results on multiple datasets demonstrate the superiority of KSCU. For instance, in the NSFW unlearning experiment, KSCU achieves a remarkable unlearning accuracy of 96.5% with FID of just 14.1, significantly outperforming the previous SOTAs.
2 Related Work
2.1 Text-to-Image Diffusion Models
@@
Diffusion models [17] were initially proposed as a generative framework based on the reverse denoising process to reconstruct target input images. With the introduction of advanced techniques such as score-based models [50], noise-conditioned score network [49], and denoising-based guidance [9], these models have become increasingly popular for text-to-image (T2I) generation. The emergence of latent diffusion models (LDMs) [40] and classifier-free guidance [16] has further improved the controllability and efficiency of diffusion-based T2I systems. Their ability to generate high-quality, semantically coherent images has established diffusion models as the dominant approach in T2I generation.
Building on these advancements, recent research has focused on optimizing T2I diffusion models in terms of efficiency, scalability, and applicability. Significant breakthroughs have been made in acceleration algorithms [48, 25, 28, 29], model architectures [40], and large-scale datasets [45, 4, 37, 38], all of which have contributed to the widespread adoption of these models in practical applications [38, 31, 42, 1]. However, despite their advantages, these models are vulnerable to misuse, leading to concerns over NSFW content [39], unauthorized depictions of individuals, and copyright-infringing artwork [1]. As T2I diffusion models continue to advance in realism and fidelity, addressing their ethical and legal challenges has become an urgent priority for the research community.
@@
Figure 2: The KSCU framework consists of three key modules: Key Step Unlearning Loss, Key Step Table, and Prompt Augmentation. The Key Step Table predefines step selection strategies for four concept unlearning tasks (Class, Style, NSFW, and Instance). During training, a denoising step $t$ is selected from the Key Step Table. The unlearn model performs denoising sampling to obtain $x_{t}$, while the Prompt Augmentation module converts the target concept $c$ into an augmented prompt $c^{*}$. Using $x_{t}$, $t$, $c$, and $c^{*}$, the Key Step Unlearning Loss is computed via Equation (7), and designated parameters of the unlearn model are updated through backpropagation.
2.2 Diffusion Model Concept Unlearning
@@
Large-scale datasets [45, 4, 37, 38] often contain undesirable data, including content with inherent biases, ethically problematic information, or copyrighted materials that may later become restricted. Models trained on such datasets acquire the ability to generate this sensitive content, posing significant security and ethical risks in text-to-image (T2I) generation. To mitigate these concerns, concept unlearning in diffusion models has been proposed, aiming to selectively erase specific generative capabilities while preserving the model’s ability to generate unrelated content—without requiring full retraining.
Several existing methods have been developed for concept unlearning in T2I diffusion models. ESD [12] and CA [23] employ a frozen guiding model to provide negative guidance for fine-tuning. UCE [13] and RECE [14] leverages a closed-form solution to modify cross-attention weights, enabling simultaneous unlearning of multiple concepts. Salun [11] and Scissorhands [52] optimize parameter updates based on their influence on concept retention. Meanwhile, RACE [21] and AdvUnlearn [59]generate adversarial samples to enhance unlearning and improve model robustness.
Although concept unlearning methods provide a certain level of security for T2I diffusion models, their effectiveness remains questionable when faced with diverse real-world scenarios and adversarial examples. Recent studies on adversarial attacks [51, 6, 58] against T2I DMs indicate that existing concept unlearning methods remain vulnerable to such attacks. To enhance robustness, current approaches rely on increasing the number of fine-tuning iterations or incorporating adversarial training [59, 21]. However, these methods significantly increase computational cost while further degrading the model’s generative quality.
Unlike previous approaches [23, 12, 53, 33], KSCU explicitly considers the relationship between the stepwise denoising process of diffusion models and the effectiveness of concept unlearning. Instead of modifying the model across all denoising steps, KSCU selectively applies unlearning to specific denoising steps that contribute most significantly to the target concept’s generation. This targeted approach reduces the number of training iterations required for unlearning, thereby minimizing unnecessary parameter adjustments and preserving the model’s overall generative quality.
3 Key Step Concept Unlearning
@@
As shown in Figure 2, our framework consists of three main components: the Key Step Unlearning Loss module, including the unlearn model that undergoes fine-tuning and the guidance model that provides reverse guidance; the Key Step Table module, which controls the denoising steps during training; and the Prompt Augmentation module to help the model learn more generalized semantic distribution of the unlearned concept in the text embedding space.
3.1 Background
Diffusion models [17] are a class of generative models that learn data distributions by gradually denoising a normally distributed variable. The process consists of two phases: the forward diffusion process and the reverse denoising process. In the forward process, Gaussian noise is incrementally added to the data over a series of time steps, transforming the data into a nearly isotropic Gaussian distribution. Conversely, the reverse process learns to iteratively denoise the data, recovering the original distribution.
Latent Diffusion Models. While traditional diffusion models operate directly in the high-dimensional data space, latent diffusion models (LDMs) [40] introduce a more efficient approach by performing the diffusion process in a compressed latent space. LDMs first employ an autoencoder to encode the input data $\mathbf{x}$ into a lower-dimensional latent representation $\mathbf{z}=E(\mathbf{x})$, where $E$ is the encoder. The diffusion process is then applied to $\mathbf{z}$ instead of $\mathbf{x}$, significantly reducing computational complexity.
The key distinction of LDMs lies in their separation of the diffusion process from the high-dimensional data space. The forward process in the latent space is defined as:
$q(\mathbf{z}_{t}|\mathbf{z}_{t-1})=\mathcal{N}(\mathbf{z}_{t};\sqrt{1-\beta_{t}}\mathbf{z}_{t-1},\beta_{t}\mathbf{I}),$ (1)
where $\mathbf{z}_{t}$ is the latent representation at time step $t$. The reverse process learns to denoise the latent representation:
$p_{\theta}(\mathbf{z}_{t-1}|\mathbf{z}_{t})=\mathcal{N}(\mathbf{z}_{t-1};\mu_{\theta}(\mathbf{z}_{t},t),\Sigma_{\theta}(\mathbf{z}_{t},t)).$ (2)
After the reverse process, the decoder $D$ maps the denoised latent representation $\mathbf{z}_{0}$ back to the data space: $\mathbf{x}=D(\mathbf{z}_{0})$. This two-stage approach not only reduces computational costs but also enables the generation of high-resolution samples by focusing on a compact latent space.
Classifier-Free Guidance. Classifier-Free Guidance [16] is a technique designed to improve the controllability of diffusion models without relying on an external classifier. Instead of conditioning the model on explicit labels, classifier-free guidance jointly trains a conditional and an unconditional diffusion model. Formally, let $\epsilon_{\theta}(\mathbf{x}_{t},t,\mathbf{c})$ denote the conditional prediction and $\epsilon_{\theta}(\mathbf{x}_{t},t)$ denote the unconditional prediction, where $\mathbf{c}$ is the conditioning information (e.g., a class label or text embedding). The guided prediction is computed as:
$\hat{\epsilon}_{\theta}(\mathbf{x}_{t},t,\mathbf{c})=\epsilon_{\theta}(\mathbf{x}_{t},t)+w\cdot(\epsilon_{\theta}(\mathbf{x}_{t},t,\mathbf{c})-\epsilon_{\theta}(\mathbf{x}_{t},t))$ (3)
where $w$ is the guidance scale controlling the strength of conditioning. A higher $w$ emphasizes the conditioning information, while $w=0$ reduces the model to an unconditional generator.
3.2 Key Step Unlearning Loss
Figure 3: Sampling with ”nudity woman” prompt reveals emergence of nudity concepts in the final 10 denoising steps.
According to formula (3), in a classifier-free guidance diffusion model, the linear extrapolation of $\epsilon_{\theta}(\mathbf{x}_{t},t,\mathbf{c})$ and $\epsilon_{\theta}(\mathbf{x}_{t},t)$ serves as the gradient of the implicit classifier $\nabla_{x_{t}}\log p(y|x_{t})$. Our objective is not to generate images corresponding to the text prompt $c$, but rather to shift the model’s predicted noise $\hat{\epsilon}_{\theta}(\mathbf{x}_{t},t,\mathbf{c})$ in the opposite direction of $\nabla_{x_{t}}\log p(y|x_{t})$. Furthermore, we exclusively fine-tune the conditional noise $\hat{\epsilon}_{\theta}(\mathbf{x}_{t},t,\mathbf{c})$ to ensure that the generation of unrelated concepts remains unaffected. The fine-tuning objective is formulated as:
$\begin{split}\hat{\epsilon}^{*}_{\theta}(\mathbf{x}_{t},t,\mathbf{c})&=\epsilon_{\theta}(\mathbf{x}_{t},t)-w\cdot(\epsilon_{\theta}(\mathbf{x}_{t},t,\mathbf{c})-\epsilon_{\theta}(\mathbf{x}_{t},t))\\ &=\epsilon_{\theta}(\mathbf{x}_{t},t)+w\cdot(\epsilon_{\theta}^{*}(\mathbf{x}_{t},t,\mathbf{c})-\epsilon_{\theta}(\mathbf{x}_{t},t))\end{split}$ (4)
From this, we derive the unlearning loss:
$\begin{split}\text{Loss}_{\text{unlearn}}&=\left\|\epsilon_{\theta}^{*}(\mathbf{x}_{t},t,\mathbf{c})-(\epsilon_{\theta}(\mathbf{x}_{t},t)-\left(\epsilon_{\theta}(\mathbf{x}_{t},t,\mathbf{c})-\epsilon_{\theta}(\mathbf{x}_{t},t)\right))\right\|_{2}\\ \end{split}$ (5)
where $\epsilon^{*}_{\theta}$ represents the unlearned model, $\epsilon_{\theta}$ is the guidance model, $x_{t}$ denotes the denoised image at step $t$, and $c$ corresponds to the text prompt.
More importantly, the Key Step Unlearning Loss also accounts for the optimization of unconditional noise. To preserve the model’s ability to generate unrelated concepts, we introduce a regularization term based on unconditional noise prediction to minimize parameter changes. Specifically, we constrain the L2 distance between the unconditional noise predictions of the guidance and unlearned models, offering partial protection for shared parameters. Additionally, we observe that unconditional noise, which uses the previous sampling step as a prior, gradually aligns with the target concept as denoising progresses. As shown in Figure 3, we use the prompt ”nudity woman” with DDIM steps set to 50 and sample images using Stable Diffusion v1-4. From left to right, the images correspond to step = 40, step = 50, and the last 10 steps sampled using only unconditional noise. The qualitative results show that even unconditional noise alone causes the generation to drift toward the ”nudity” concept. This highlights the necessity of applying unlearning not only to the conditional branch but also to the unconditional branch in classifier-free guidance diffusion models. Accordingly, we apply unlearning training to the unconditional component as well. The regularization loss is formulated as:
$\text{Loss}_{\text{regular}}=\left\|\epsilon_{\theta}^{*}(\mathbf{x}_{t},t)-((1+\lambda_{1})\epsilon_{\theta}(\mathbf{x}_{t},t)-\lambda_{1}\ \epsilon_{\theta}(\mathbf{x}_{t},t,\mathbf{c}))\right\|^{2}_{2}$ (6)
where $\lambda_{1}$ needs to be a small value that is linearly related to $t$. For $t$ from 1000 to 0, we set $\lambda_{1}=(1000-t)/(2*10^{5})$
The final Key Step Unlearning Loss function for KSCU is defined as:
where $\lambda_{2}=1\times 10^{-4}$ in KSCU.

### v3

@@ Abstract
Text-to-image diffusion models (T2I DMs), represented by Stable Diffusion, which generate highly realistic images based on textual input, have been widely used, but their flexibility also makes them prone to misuse for producing harmful or unsafe content. Concept unlearning has been used to prevent text-to-image diffusion models from being misused to generate undesirable visual content. However, existing methods struggle to trade off unlearning effectiveness with the preservation of generation quality. To address this limitation, we propose Key Step Concept Unlearning (KSCU), which selectively fine-tunes the model at key steps to the target concept. KSCU is inspired by the fact that different diffusion denoising steps contribute unequally to the final generation. Compared to previous approaches, which treat all denoising steps uniformly, KSCU avoids over-optimization of unnecessary steps for higher effectiveness and reduces the number of parameter updates for higher efficiency. For example, on the I2P dataset, KSCU outperforms ESD by 8.3% in nudity unlearning accuracy while improving FID by 8.4%, and achieves a high overall score of 0.92, substantially surpassing all other SOTA methods.
I Introduction
In recent years, the advancement of text-to-image (T2I) diffusion models [1, 2, 3, 4, 5, 6] has led to their widespread adoption across various domains [7, 8, 9, 10, 11], including short videos, comics, and illustrations. These models enable users to translate natural language prompts into high-quality, semantically coherent visual outputs, greatly lowering the entry barrier for creative applications. As a result, AI-generated images have become an integral part of daily life for many individuals. However, large-scale T2I models are trained on extensive datasets [12, 13, 14, 7] that inevitably contain sensitive or problematic content, such as unauthorized artistic works [15, 16], culturally biased representations, or Not Safe For Work (NSFW) material. Consequently, these models can inadvertently regenerate infringing or inappropriate content when prompted with relevant text inputs [17]. This capability raises pressing concerns regarding copyright infringement, ethical use, and social security risks, highlighting the urgent need for effective mitigation strategies.
@@ I Introduction
The generation of NSFW content poses a major security risk for mainstream text-to-image diffusion models. To mitigate this issue, researchers have explored various approaches, including dataset-based screening, post-hoc content filtering [18], and fine-tuning techniques [19, 20, 21]. While dataset curation attempts to remove harmful samples at the source, it is costly and incomplete, as sensitive content is often interwoven with valuable training data. Filtering mechanisms at inference time can suppress unsafe outputs but fail to prevent the model from retaining harmful capabilities internally. By contrast, concept unlearning has emerged as a promising fine-tuning-based solution [22, 23, 24, 25, 26, 27, 28], which selectively adjusts partial model parameters to remove the model’s ability to generate specific concepts. Through concept unlearning, T2I diffusion models effectively ‘forget’ designated visual concepts by replacing them with semantically neutral or degraded alternatives. This targeted parameter adjustment (also known as concept erasure [29, 30]) offers distinct advantages over conventional approaches—providing greater robustness than post-hoc filtering [18] while avoiding the intensive labor of dataset curation [31]. Consequently, concept unlearning has become one of the most promising strategies to curb the generation of NSFW content in diffusion models.
Although concept unlearning for T2I diffusion models has made notable progress, key challenges remain. (a) Blind unlearning across all diffusion steps. Existing methods often overlook the stepwise nature of denoising in diffusion models. The generation of visual concepts is hierarchical: early steps capture coarse structures and low-frequency semantics, while later steps refine fine-grained textures and high-frequency details [32, 33, 34, 35, 36]. Ignoring this temporal structure leads to unnecessary parameter updates and reduced efficiency. We hypothesize and empirically validate that effective concept unlearning can be achieved by selectively modifying only a small set of key steps that significantly influence the final output. (b) Unbalanced unlearning effectiveness and generative retainability. Existing approaches often face a trade-off: aggressive parameter updates can achieve stronger forgetting but degrade the generation of unrelated concepts (over-unlearning), while conservative updates preserve retainability but result in incomplete removal of the target concept. Achieving a principled balance between these two objectives remains an open challenge.
@@
To address these challenges, we propose Key Step Concept Unlearning (KSCU), a novel framework for concept unlearning in T2I diffusion models. Unlike prior methods that operate across the entire denoising trajectory, KSCU selectively fine-tunes only a subset of key steps determined by a predefined Key Step Table, as illustrated in Figure 1. For different unlearning tasks, multiple step intervals are empirically defined, and a step-selection algorithm prioritizes those with the greatest influence on the final output, enabling efficient and targeted unlearning. To further enhance robustness, KSCU applies Prompt Augmentation in the final training stage, where augmented prompts help the model better capture the semantic distribution of the unlearned concept in the embedding space. Additionally, we introduce a Key Step Unlearning Optimization tailored to Classifier-Free Guidance (CFG) diffusion models, ensuring that the erasure process interacts smoothly with guidance scaling. By combining Key Step Table, Prompt Augmentation, and Key Step Unlearning Optimization, KSCU achieves more effective concept unlearning while better preserving generative retainability.
@@
  • We propose KSCU, a novel concept unlearning framework for T2I diffusion models. Unlike prior works, we explicitly observe the hierarchical concept formation process in diffusion and demonstrate that unlearning can be effectively achieved by updating only a small set of key steps. KSCU fine-tunes the model exclusively at these steps with minimal adjustments, enabling efficient erasure while preserving generative retainability.
@@
  • KSCU integrates three key innovations: Key Step Table, a step-selection strategy that focuses on critical denoising steps to achieve effective unlearning with minimal overhead; Prompt Augmentation, which enhances robustness by enriching semantic coverage through diverse prompt variations; and Key Step Unlearning Optimization, an optimization scheme tailored for CFG-based diffusion models, enabling effective concept removal while maintaining high image quality.
  • We conduct extensive experiments on multiple benchmarks, covering NSFW, style, and category unlearning. Results show that KSCU consistently outperforms state-of-the-art baselines. For instance, in NSFW unlearning, KSCU achieves an overall score of 0.92 and an unlearning accuracy of 96.5% with an FID of 14.1, significantly surpassing previous SOTAs.
II Background & Related Work
II-A Text-to-Image Diffusion Models
@@
Diffusion models [37] were initially proposed as a generative framework based on the reverse denoising process to reconstruct target input images. Compared with earlier paradigms such as GANs [38] and VAEs [39], diffusion-based approaches demonstrate superior training stability, more faithful mode coverage, and better scalability to high-dimensional distributions. With the introduction of advanced techniques such as score-based models [40], noise-conditioned score networks [41], and denoising-based guidance [3], these models have become increasingly popular for text-to-image (T2I) generation.
The emergence of Latent Diffusion Models (LDMs) [31] has further improved scalability and efficiency. By operating in a learned latent space, LDMs substantially reduce the computational burden while retaining high generative fidelity, making them feasible for large-scale training. Specifically, given an image $\mathbf{x}$, an encoder $E(\cdot)$ maps it to the latent code $\mathbf{z}=E(\mathbf{x})$.
The forward diffusion process then gradually corrupts the latent with Gaussian noise over $T$ steps:
$\mathbf{z}_{t}=\sqrt{\bar{\alpha}_{t}}\,\mathbf{z}_{0}+\sqrt{1-\bar{\alpha}_{t}}\,\boldsymbol{\epsilon},\quad\boldsymbol{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I}),$ (1)
where $\bar{\alpha}_{t}=\prod_{i=1}^{t}(1-\beta_{i})$ is the cumulative product of noise schedule coefficients.
During inference, the model reverses this process by predicting the added noise $\boldsymbol{\epsilon}$ at each step and updating the latent accordingly:
$\mathbf{z}_{t-1}=\frac{1}{\sqrt{\alpha_{t}}}\left(\mathbf{z}_{t}-\frac{1-\alpha_{t}}{\sqrt{1-\bar{\alpha}_{t}}}\,\hat{\boldsymbol{\epsilon}}_{\theta}\right)+\sigma_{t}\,\boldsymbol{\eta},\quad\boldsymbol{\eta}\sim\mathcal{N}(\mathbf{0},\mathbf{I}),$ (2)
where $\hat{\boldsymbol{\epsilon}}_{\theta}$ is the predicted noise, and $\sigma_{t}$ is typically set according to the scheduler.
To enhance controllability, Classifier-Free Guidance (CFG) [4] interpolates between conditional and unconditional predictions:
$\hat{\boldsymbol{\epsilon}}_{\theta}=\boldsymbol{\epsilon}_{\theta}(\mathbf{z}_{t},t)+w\cdot\left(\boldsymbol{\epsilon}_{\theta}(\mathbf{z}_{t},t,\mathbf{c})-\boldsymbol{\epsilon}_{\theta}(\mathbf{z}_{t},t)\right),$ (3)
where $w$ denotes the guidance scale, and $\boldsymbol{\epsilon}_{\theta}(\cdot)$ and $\boldsymbol{\epsilon}_{\theta}(\cdot,\mathbf{c})$ represent the unconditional and conditional predictions, respectively.
Together, LDMs and CFG have established diffusion models as the dominant approach in T2I generation. Their ability to generate high-quality, semantically coherent images, even under complex compositional prompts, has set a new benchmark in the field.
On this basis, recent research has focused on optimizing T2I diffusion models in terms of efficiency, scalability, and applicability. Significant breakthroughs have been made in acceleration algorithms [42, 43, 44, 45], model architectures [31], and large-scale datasets [12, 13, 14, 7], all of which have contributed to the widespread adoption of these models in practical applications [7, 8, 9, 10]. Beyond visual quality, diffusion models have also demonstrated strong generalization in multimodal and cross-domain settings, such as image editing, 3D synthesis, and video generation [46, 47].
However, despite their advantages, these models are vulnerable to misuse, leading to concerns over NSFW content [18], unauthorized depictions of individuals, and copyright-infringing artwork [10]. Furthermore, the sheer expressive power of diffusion models introduces unique risks, such as prompt injection attacks, data leakage, and ethical dilemmas in human–AI collaboration. As T2I diffusion models continue to advance in realism and fidelity, addressing their ethical and legal challenges has become an urgent priority for the research community. Efforts in controllable generation, safety-aware fine-tuning, and concept erasure are increasingly viewed as essential to ensure their responsible deployment.
@@
Figure 2: The KSCU framework consists of three key modules: Key Step Table, Prompt Augmentation, and Key Step Unlearning Optimization. During training, a denoising step $t$ is selected from the Key Step Table. The unlearn model performs denoising sampling to obtain $z_{t}$, while the Prompt Augmentation module converts the target concept $c$ into an augmented prompt $c^{*}$. Using $z_{t}$, $t$, $c$, and $c^{*}$, the Key Step Unlearning Optimization is computed via Equation (7), and designated parameters of the unlearn model are updated through backpropagation.
II-B Diffusion Model Concept Unlearning
@@
Large-scale datasets [12, 13, 14, 7] often contain undesirable data, including content with inherent biases, ethically problematic information, or copyrighted materials that may later become restricted. When trained on such datasets, T2I diffusion models inevitably acquire the ability to reproduce these sensitive patterns, raising significant security and ethical concerns. To address this issue, concept unlearning has been introduced as a promising direction, aiming to selectively erase specific generative capabilities while preserving the model’s ability to synthesize unrelated content—without requiring costly full retraining.
A variety of approaches have been proposed to realize this goal. Early methods such as ESD [29] and CA [22] rely on a frozen guiding model to provide negative supervision during fine-tuning, effectively discouraging the generation of undesired concepts while leaving unrelated knowledge largely intact. UCE [25] and RECE [26] instead adopt closed-form updates to cross-attention weights, enabling efficient unlearning of multiple concepts simultaneously. Parameter-significance-based approaches like Salun [23] and Scissorhands [48] update weights selectively according to their contribution to concept retention, minimizing interference with other learned capabilities. In parallel, adversarial strategies such as RACE [30] and AdvUnlearn [49] generate challenging counterexamples, which improve both the effectiveness and robustness of unlearning. More recently, SPEED [50] and MACE [51] have advanced scalable multi-concept unlearning: SPEED formulates erasure as a null-space constrained optimization for efficient removal while preserving unrelated priors, whereas MACE leverages cross-attention refinement and modular LoRA fusion to support large-scale erasure across hundreds of concepts.
Despite these advances, existing methods provide only partial mitigation. Their reliability remains limited when confronted with diverse real-world scenarios or adversarial prompts. Recent studies on attacks against T2I diffusion models [52, 53, 54] further highlight that current unlearning techniques are vulnerable, underscoring the need for more effective and robust solutions.
III Key Step Concept Unlearning
@@
As shown in Figure 2, our framework consists of three main components: the Key Step Table module, which controls the denoising steps during training; the Prompt Augmentation module to help the model learn a more generalized semantic distribution of the unlearned concept in the text embedding space; and the Key Step Unlearning Optimization module, including the unlearned model that undergoes fine-tuning and the guidance model that provides reverse guidance.
III-A Motivation & Key Step Table Generation
Figure 3: Correlation between text and attention map at different steps
TABLE I: Preliminary Verification Experiment. The superior UA and UDA performance of ESD early 70% steps is attributed to model collapse. While ESD last 70% steps shows comparable unlearning effectiveness to full-step ESD, with less damage to the generative capability and improved time efficiency.
[…]
ESD 88.2 73.7 15.4 2557
ESD last 70% steps 85.9 80.9 15.2 2072
ESD early 70% steps 90.4 57.7 69.7 1355
Due to the stepwise nature of diffusion-based generation, each sampling step contributes differently to the final image. Existing work indicates that earlier denoising steps predominantly capture low-frequency structures, while later steps progressively refine high-frequency details [33, 32]. This leads to a layered emergence of semantic concepts during generation, as illustrated in Figure 3. Motivated by this observation, we hypothesize that concept unlearning can be achieved by modifying only the later denoising steps to adjust high-frequency information with greater semantic impact, replacing the target concept with a nearby but semantically distinct alternative (e.g., wolf $\rightarrow$ dog), thereby preventing the unwanted concept from appearing in the generated output.
To validate this hypothesis, we conduct a preliminary experiment using the ESD method [29] on Stable Diffusion v1-4, restricting optimization to either the first or last 70% of denoising steps. As shown in Table I, training on the early 70% leads to model collapse, while fine-tuning the last 70%—with 30% fewer iterations—retains 95% of the full-step unlearning accuracy. These results indicate that targeting high-frequency denoising steps has the potential to achieve effective concept unlearning while better preserving the model’s generative capability with improved time efficiency.
Based on the above conclusion, we design a step selection algorithm to generate Key Step Table. The Key Step Table consists of incrementally growing step intervals and shifts the starting point forward after a fixed number of loops to progressively cover a broader range of late denoising steps. As shown in Algorithm 1, the algorithm constructs the Key Step Table from a given starting step $S$ to an ending step $E$. In each iteration, the full interval from $s_{\text{cur}}$ to $E$ is appended to the Key Step Table. After $loop_{n}$ iterations, the starting step $s_{\text{cur}}$ is shifted forward by one, and the process is repeated until the Key Step Table reaches the desired length $L$. This mechanism ensures that training focuses more on later denoising steps, which are more influential, while avoiding overfitting to a single segment.